### v1

# On Araki-Type Trace Inequalities

Abstract.
In this paper, we prove a trace inequality $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\big{[}f(A)(A^{1/2}BA^{1/2})^{s}\big{]}$ for any positive and monotone increasing function $f$, $s\in[0,1]$, and positive semi-definite matrices $A$ and $B$. On the other hand, for $s\in[0,1]$ such that the map $x\mapsto x^{s}g(x)$ is positive and decreasing, then $\operatorname{Tr}\big{[}g(A)(A^{1/2}BA^{1/2})^{s}\big{]}\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right]$.
1. Introduction
Golden–Thompson trace inequality [1, 2, 3], i.e., for Hermitian matrices $H,K$,
$\operatorname{Tr}\left[\mathrm{e}^{H+K}\right]\leq\operatorname{Tr}\left[\mathrm{e}^{H}\mathrm{e}^{K}\right],$ (1)
and the Araki–Lieb–Thirrling inequality [4, 5]: for $A,B\geq 0$,
$\operatorname{Tr}\left[\left(A^{1/2}BA^{1/2}\right)^{r}\right]\leq\operatorname{Tr}\left[A^{r/2}B^{r}A^{r/2}\right],\quad r\geq 1$ (2)
(the inequality reverses [^1] In some literature, (2) for $r\geq 1$ is called Lieb–Thirrling inequality [4], and the reversed inequality for $r\in[0,1]$ is sometimes called the Araki–Cordes inequality [5, 6]. In this paper, we will use the term, Araki–Lieb–Thirrling inequality, referring to the inequality for all $r\geq 0$. for $r\in[0,1]$) plays important roles in statistical mechanics, quantum information theory, and quantum computing. For example, (1) implies that
$D_{\alpha}(\rho\|\sigma):=\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\rho^{\alpha}\sigma^{1-\alpha}\right]\leq\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\operatorname{\mathrm{e}}^{\alpha\log\rho+(1-\alpha)\log\sigma}\right]=:D_{\alpha}^{\flat}(\rho\|\sigma),\quad\forall\alpha\in(0,1),$ (3)
where $D_{\alpha}(\rho\|\sigma)$ is the Petz–Rényi divergence [7], $D_{\alpha}^{\flat}(\rho\|\sigma)$ is the log-Euclidean Rényi divergence, and $\rho,\sigma$ are density matrices. On the other hand, (2) implies that
$D_{\alpha}^{*}(\rho\|\sigma):=\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\left(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}\right)^{\alpha}\right]\leq D_{\alpha}(\rho\|\sigma),\quad\forall\alpha>0,$ (4)
where $D_{\alpha}^{*}(\rho\|\sigma)$ is the sandwiched Rényi divergence [8, 9]. Essentially, Golden–Thompson inequality and Araki–Lieb–Thirrling inequality manifest the noncommutative nature of quantum mechanics and that certain inequalities instead of the equality will be induced in analytical derivations.
Araki–Lieb–Thirrling trace inequality (2) can be strengthened to the log-majorization relation [5]:
$\left(A^{1/2}BA^{1/2}\right)^{r}\prec_{\log}A^{r/2}B^{r}A^{r/2},\quad r\geq 1$ (5)
(the log-majorization relation reverses for $r\in[0,1]$). Here, for matrices $A,B\in\mathcal{M}_{n}(\mathds{C})$ with nonnegative eigenvalues, the log-majorization of $A$ by $B$, denoted by $A\prec_{\log}B$, means that
$\prod_{i=1}^{k}\lambda_{i}(A)\leq\prod_{i=1}^{k}\lambda_{i}(B),\quad k=1,2\ldots,n,$ (6)
and $\det{A}=\det{B}$, where $\lambda_{i}(A)$ is is $i$-th eigenvalue of $A$ arranged in non-increasing order. Likewise, (1) can be strengthened by (5) and the Lie–Trotter formula [10]:
$\operatorname{\mathrm{e}}^{H+K}\prec_{\log}\left(\operatorname{\mathrm{e}}^{pH/2}\operatorname{\mathrm{e}}^{pK}\operatorname{\mathrm{e}}^{pH/2}\right)^{1/p},\quad\forall p>0.$ (7)
Ando and Hiai later obtained a log-majorization via Löwner–Heinz operator inequality and the antisymmetric tensor power trick (see e.g., [11, Lemma 2.4]), which complements Golden–Thompson type: [10]:
$\left[\mathrm{e}^{pK/2}\left(\mathrm{e}^{-pK/2}\mathrm{e}^{pH}\mathrm{e}^{-pK/2}\right)^{\alpha}\mathrm{e}^{pK/2}\right]^{1/p}\prec_{\log}\operatorname{\mathrm{e}}^{\alpha H+(1-\alpha)K},\quad\alpha\in[0,1],\;p>0.$ (8)
The above then implies that
$D_{\alpha}^{\flat}(\rho\|\sigma)\leq\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\rho\left(\rho^{-1/2}\sigma\rho^{-1/2}\right)^{1-\alpha}\right]=:\widehat{D}_{\alpha}(\rho\|\sigma),\quad\alpha\in(0,1].$ (9)
Bebiano, Lemos and Providência applied Ando and Hiai’s proof technique with Furuta’s operator inequality [12] to obtain the following log-majorization of Araki’s type.
Theorem BLP (BLP Inequality [13]).
For $A,B\geq 0$,
$A^{1+q}B^{q}\prec_{\log}A\left(A^{r/2}B^{r}A^{r/2}\right)^{q/r},\quad 0<q\leq r.$ (10)
Equivalently,
$A^{t+s}B^{s}\prec_{\log}A^{t}\left(A^{1/2}BA^{1/2}\right)^{s},\quad s\in(0,1],\;t>0.$ (11)
Theorem BLP provides an alternative proof to the following trace inequality
$\operatorname{Tr}\left[A\left(\log A+\log B\right)\right]\leq\frac{1}{p}\operatorname{Tr}\left[A\log\left(A^{p/2}B^{p}A^{p/2}\right)\right],\quad p>0$ (12)
(see also the early paper by Hiai and Petz [14], the complementaion by Ando and Hiai [10], and the recent developments by Lieb and Carlen [15]). Inequality (12) then implies that the following relation between Umegaki’s quantum relative entropy $D(\rho\|\sigma)$ [16] and Belavkin–Staszewski’s relative entropy $\widehat{D}(\rho\|\sigma)$ [17], i.e.,
$D(\rho\|\sigma):=\operatorname{Tr}\left[\rho(\log\rho-\log\sigma)\right]\leq\operatorname{Tr}\left[\rho\log\left(\rho^{1/2}\sigma^{-1}\rho^{1/2}\right)\right]=:\widehat{D}(\rho\|\sigma).$ (13)
With efforts to develop the reversed BLP inequalities [18, 19, 20], the above log-majorizations can be unified as the following generalized BLP (GBLP) inequalities.222 Some literature studied different generalizations of the BLP inequality, (11), and used the term generalized BLP [21, 22, 23, 24]. In this paper, we call the inequalities and the reversed versions with the possible range of parameters as in Theorem GBLP as the GBLP inequality.
Theorem GBLP (GBLP Inequality [13, 18, 19, 20]).
For $A,B\geq 0$,
$A^{r+q}B^{q}\prec_{\log}A^{r}\left(A^{p/2}B^{p}A^{p/2}\right)^{q/p},\quad 0<q\leq p,\;r\geq 0.$ (14)
If either $0\leq r\leq p\leq q$ and $p>0$, or $0\leq q<p$ and $-r\geq q$, then (14) holds with reversed log-majorization.
Theorem GBLP and together with a result by Matharu and Aujla (obtained similarly from the Furuta inequality) [25] show the following relations:
$\begin{split}A\left(A^{-1/2}BA^{-1/2}\right)^{s}&\preceq_{\log}A^{1-s}B^{s},\quad s\in[0,1]\cup[2,\infty);\\ A\left(A^{-1/2}BA^{-1/2}\right)^{s}&\succeq_{\log}A^{1-s}B^{s},\quad s\in[1,2].\end{split}$ (15)
The relations given in (15) is intriguing in the following sense. Considering $s\in[0,1]$, then Araki–Lieb–Thirrling inequality shows that $\left(A^{-1/2}BA^{-1/2}\right)^{s}\succeq_{\log}A^{-s}B^{s}$. However, when multiplying both sides by $A$, distributing the power $s$ yields a reversed log-majorization $A\left(A^{-1/2}BA^{-1/2}\right)^{s}\preceq_{\log}A^{1-s}B^{s}$. Hence, the goal of this paper is to identify which function $f$ such that the following trace inequality holds,
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],$ (16)
or when the trace inequality reverses. Obviously, the BLP inequality (11) already provides an answer to (16) for $f(A)=A^{t}$, $t>0$. However, whenever $f$ is not a power function, the core idea of proving Theorems BLP and GBLP, i.e., the antisymmetric tensor power trick with the Furuta inequality, is not applicable.
Our main result is to show that any nonnegative and nondecreasing function $f$ satisfies (16) (see Theorem 4), and, for any $s\in[0,1]$ and any function $g$ such that $x\mapsto x^{s}g(x)$ is nonnegative and nonincreasing, (16) is reversed (see Proposition 5).
We provide our main results and the proofs in Section 2. In Section 3, we discuss the scneario for $s>1$.
2. An Araki-type trace inequality
We first show that (16) holds if $f(A)$ is a kind of orthogonal projection matrices.
Proposition 1.
Let $A=\sum_{i=1}^{n}\lambda_{i}P_{i}$ be the spectral decomposition of positive semi-definite $A$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, and let $\Pi_{k}:=\sum_{i=1}^{k}P_{i}$. Then, for all $B\geq 0$,
[…]
Remark 2.1.
The proposition 1 holds for $\Pi_{k}$ being a projection onto the eigenspace associated with the $k$-th largest eigenvalues. The trace inequality may not be true once replacing $\Pi_{k}$ by an arbitrary eigenprojection $P_{i}$. To see a counter-example, let $s=1/2$ and
$A=\begin{bmatrix}2&0\\ 0&1\end{bmatrix},\,B=\begin{bmatrix}1&1\\ 1&1\end{bmatrix},\,P_{2}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}.$ (17)
Then, $\operatorname{Tr}[P_{2}A^{s}B^{s}]=1/\sqrt{2}\not\leq\operatorname{Tr}[P_{2}(A^{1/2}BA^{1/2})^{s}]\approx 0.5774$.
In the following proofs and lemmas, we suppose $A>0$, and the case $A\geq 0$ then follows by taking the limit. To prove Proposition 1, we need the following two lemmas.
Lemma 2.
Following the notation in Proposition 1, we denote $I-\Pi_{k}$ by $\Pi_{k}^{\text{c}}$, and let $\tilde{A}_{k}=\Pi_{k}A+\lambda_{k}\Pi_{k}^{\textnormal{c}}=\sum_{i=1}^{n}\max\{\lambda_{i},\lambda_{k}\}P_{i}$. Then, we have
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
The inequality reverses for all $s\in[1,2]$.
Proof.
First, note that $A$ and each $\tilde{A}_{k}$ commute. Let $R=(A\tilde{A}_{k}^{-1})^{1/2}$ be a Hermitian contraction, i.e., $R^{2}\leq I$. Since $x\mapsto x^{s}$ is operator concave for $s\in[0,1]$, the operator Jensen’s inequality [26] implies that
$\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}=R\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}R\leq\left(R\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}R\right)^{s}=\left(A^{1/2}BA^{1/2}\right)^{s}.$
Hence, we obtain
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]$ $=\operatorname{Tr}\left[\Pi_{k}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\right]$ (18)
$\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],$ (19)
with the fact that $\Pi_{k}(A\tilde{A}_{k}^{-1})^{1/2}=(A\tilde{A}_{k}^{-1})^{1/2}\Pi_{k}=\Pi_{k}$.
The case of $s\in[1,2]$ follows similarly by the operator convexity of $x\mapsto x^{s}$, $s\in[1,2]$. ∎
Note that the BLP inequality implies that, for $A,B\geq 0$,
$\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right]\leq\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],\,t\geq 0.$ (20)
We have another lemma.
Lemma 3.
For $A>0$ and $B\geq 0$,
$\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right],\quad\forall\,s\in[0,1],\,t\leq-s.$
Proof.
Substitute $B$ by $A^{-1/2}BA^{-1/2}$ to (20). Then
$\operatorname{Tr}\left[A^{t}A^{s}\left(A^{-1/2}BA^{-1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}B^{s}\right],\quad\forall\,s\in[0,1],\,t\geq 0.$
Next, substitute $A$ by $A^{-1}$, then
$\operatorname{Tr}\left[A^{-t-s}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{-t}B^{s}\right],\quad\forall\,s\in[0,1],\;t\geq 0.$
Or,
$\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right],\quad\forall\,s\in[0,1],\;t\leq-s.$
Proof of Proposition 1.
Combining the two different conditions regarding $t$ in (20) and Lemma 3, we conclude that for $f(x)=x^{t}$, $t\geq 0$ or $f(x)=-x^{t}$, $t\leq-s$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$ (21)
Notice that $\operatorname{Tr}[\Pi_{k}A^{s}B^{s}]=\operatorname{Tr}[\Pi_{k}\tilde{A}_{k}^{s}B^{s}]$ for $\Pi_{k}$ introduced in Lemma 2. Hence, to complete the proof of Proposition 1, by Lemma 2 we only need to show that
$\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$ (22)
To show this, we first introduce $\tilde{A}_{k,\varepsilon}=\Pi_{k}A+(\lambda_{k}-\varepsilon)\Pi_{k}^{\mathrm{c}}$ as an approximation of $\tilde{A}_{k}$, for $\varepsilon\in(0,\lambda_{k})$, where $\lambda_{k}$ is $k$-th largest eigenvalue of $A$.
Define
$g_{t}(x)=1-\left(\frac{x}{\lambda_{k}-\varepsilon}\right)^{t},\quad\forall\,t\leq-s,$
then, (21) implies that
$\operatorname{Tr}\left[g_{t}(\tilde{A}_{k,\varepsilon})\tilde{A}_{k,\varepsilon}^{s}B^{s}\right]\leq\operatorname{Tr}\left[g_{t}(\tilde{A}_{k,\varepsilon})\left(\tilde{A}_{k,\varepsilon}^{1/2}B\tilde{A}_{k,\varepsilon}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],\,t\leq-s,$
with noticing that the constant function $1$ is just $f(x)=x^{t}$ at $t=0$.
Let $t\to-\infty$, then $g_{t}(\tilde{A}_{k,\varepsilon})\to\Pi_{k}$, and hence
$\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k,\varepsilon}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k,\varepsilon}^{1/2}B\tilde{A}_{k,\varepsilon}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
Finally, by letting $\varepsilon\to 0$, we obtain
$\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
In conclusion, for all $s\in[0,1]$, we have
$\operatorname{Tr}\left[\Pi_{k}A^{s}B^{s}\right]=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],$
which completes the proof of Proposition 1. ∎
Now we are in a position to prove our main results.
Theorem 4.
Let $A,B\geq 0$. For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
Proof.
Since $f$ is nonnegative and nondecreasing, we can choose some appropriate nonnegative coefficients $c_{k}$ for each $k=1,\cdots,n$, such that $f(A)=\sum_{k=1}^{n}c_{k}\Pi_{k}$. Then, the claim follows from Proposition 1. ∎
As above, we may also apply substitutions to obtain another direction of the same inequality, but with a different condition.
Proposition 5.
Let $A,B\geq 0$. For any $s\in[0,1]$ and any function $g$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, such that $x\mapsto x^{s}g(x)$ is nonnegative and nonincreasing, then
$\operatorname{Tr}\left[g(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].$
Proof.
Again, we assume $A>0$, and extend to $A\geq 0$ via a limit. Set $f(x)=x^{-s}g(x^{-1})$, which is nonnegative and nondecreasing. Since the domain of $f$ includes the spectrum of $A^{-1}$, by Theorem 4,
$\operatorname{Tr}\left[g(A)\left(A^{{1}/{2}}BA^{{1}/{2}}\right)^{s}\right]$ $=\operatorname{Tr}\left[f\left(A^{-1}\right)\left(A^{-1}\right)^{s}\left(A^{{1}/{2}}BA^{{1}/{2}}\right)^{s}\right]$ (23)
$\leq\operatorname{Tr}\left[f\left(A^{-1}\right)\left(A^{-1/2}\cdot A^{{1}/{2}}BA^{{1}/{2}}\cdot A^{-1/2}\right)^{s}\right]$ (24)
$=\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].$ (25)
3. A Conjecture for $s>1$
Theorem 4 shows the trace inequality for all $s\in[0,1]$. Inspired by the reverse BLP inequality [18, 21, 22], we conjecture the following reverse direction.
Conjecture 1.
For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s>1.$
Note that $x\mapsto x^{s}$ is operator convex for $s\in[1,2]$, we may adapt the previous technique to prove Conjecture 1 for $s\in[1,2]$, provided that the condition $0\leq r\leq p\leq q$ in the reverse direction of Theorem GBLP can be relaxed to $0\leq p\leq q$. Thereby, we conjecture the following reverse log-majorization, which complements (14).
Conjecture 2.
For $A,B\geq 0$,
$\operatorname{Tr}\left[A^{r+q}B^{q}\right]\geq\operatorname{Tr}\left[A^{r}\left(A^{p/2}B^{p}A^{p/2}\right)^{q/p}\right],\quad 0\leq p\leq q,\;r\geq 0.$ (26)
(To show Conjecture 1 for $s\in[1,2]$, the trace inequality version of Conjecture 2 would be sufficient.)
In fact, we can directly prove Conjecture 1 for a special case of $s=2$ without resorting to Conjecture 2.
Proposition 6.
For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad s=2.$
Proof.
Our claim is based on Lemma 2 and the reverse direction to (22), which is presented in the following Lemma 7. ∎
Lemma 7.
Let $A=\sum_{i=1}^{n}\lambda_{i}P_{i}$ be the spectral decomposition of positive semi-definite $A$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d}\geq 0$, $\Pi_{k}:=\sum_{i=1}^{k}P_{i}$, $\Pi_{k}^{\textnormal{c}}:=I-\Pi_{k}$, and $\tilde{A}_{k}:=\Pi_{k}A+\lambda_{k}\Pi_{k}^{\textnormal{c}}$. Then,
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right],\quad s=2.$ (27)
Proof.
Using the fact that $\lambda_{k}\Pi_{k}\leq\Pi_{k}\tilde{A}_{k}\Pi_{k}$ by construction and the Araki–Lieb–Thirrling trace inequality (2), we calculate
$\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{2}\right]$ $=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{\tilde{A}_{k}}B\right]$ (28)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{\left(\Pi_{k}\tilde{A}_{k}\Pi_{k}+\lambda_{k}\Pi_{k}^{\text{c}}\right)}B\right]$ (29)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{\Pi_{k}\tilde{A}_{k}\Pi_{k}}B\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{(\lambda_{k}\Pi_{k}^{\text{c}})}B\right]$ (30)
$=\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{(\lambda_{k}\Pi_{k}^{\text{c}})}B\right]$ (31)
$=\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}{\lambda_{k}\Pi_{k}}B{\Pi_{k}^{\text{c}}}B\right]$ (32)
$\leq\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}{\Pi_{k}\tilde{A}_{k}\Pi_{k}}B{\Pi_{k}^{\text{c}}}B\right]$ (33)
$=\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (34)
(2)${\leq}\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k})^{2}(\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (35)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B\Pi_{k}B\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (36)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B(\Pi_{k}+{\Pi_{k}^{\text{c}}})B\right]$ (37)
$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}B^{2}\right].$ (38)
4. Conclusions
In this paper, we derived some Araki-type trace inequalities that involve monotone functions other than powers. Although usual Araki-type inequalities and the BLP inequalities are expressed in a stronger form of log-majorization, however, the antisymmetric tensor power trick may not be applicable for functions beyond powers. On the other hand, trace inequalities find more and immediate applications in other fields such as quantum information science and mathematical physics, which motivates us to further Araki-type trace inequalities to a broader form.
Numerical experiments suggest that Conjecture 1 for $s>1$ and a slightly general reverse BLP inequality in Conjecture 2 could be true, which is left for future work.

### v2

@@
 Abstract.
-In this paper, we prove a trace inequality $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\big{[}f(A)(A^{1/2}BA^{1/2})^{s}\big{]}$ for any positive and monotone increasing function $f$, $s\in[0,1]$, and positive semi-definite matrices $A$ and $B$.
-On the other hand, for $s\in[0,1]$ such that the map $x\mapsto x^{s}g(x)$ is positive and decreasing, then $\operatorname{Tr}\big{[}g(A)(A^{1/2}BA^{1/2})^{s}\big{]}\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right]$.
+In this paper, we prove a trace inequality $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\big[f(A)(A^{1/2}BA^{1/2})^{s}\big]$ for any positive and monotonically increasing function $f$, $s\in[0,1]$, and positive semi-definite matrices $A$ and $B$.
+On the other hand, if $s\in[0,1]$ and the map $x\mapsto x^{s}g(x)$ is positive and decreasing, then $\operatorname{Tr}\big[g(A)(A^{1/2}BA^{1/2})^{s}\big]\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right]$.
+Key words and phrases:
+Araki–Lieb–Thirring inequality, Golden–Thompson inequality, trace inequality
+1991 Mathematics Subject Classification:
+Primary 15A42, 15A45, 15A60, 47A60
 1. Introduction
-Golden–Thompson trace inequality [1, 2, 3], i.e., for Hermitian matrices $H,K$,
+The Golden–Thompson trace inequality [1, 2, 3], i.e., for Hermitian matrices $H,K$,
 $\operatorname{Tr}\left[\mathrm{e}^{H+K}\right]\leq\operatorname{Tr}\left[\mathrm{e}^{H}\mathrm{e}^{K}\right],$ (1)
-and the Araki–Lieb–Thirrling inequality [4, 5]: for $A,B\geq 0$,
+and the Araki–Lieb–Thirring inequality [4, 5]: for $A,B\geq 0$,
 $\operatorname{Tr}\left[\left(A^{1/2}BA^{1/2}\right)^{r}\right]\leq\operatorname{Tr}\left[A^{r/2}B^{r}A^{r/2}\right],\quad r\geq 1$ (2)
-(the inequality reverses [^1] In some literature, (2) for $r\geq 1$ is called Lieb–Thirrling inequality [4], and the reversed inequality for $r\in[0,1]$ is sometimes called the Araki–Cordes inequality [5, 6].
-In this paper, we will use the term, Araki–Lieb–Thirrling inequality, referring to the inequality for all $r\geq 0$.
-for $r\in[0,1]$) plays important roles in statistical mechanics, quantum information theory, and quantum computing.
+(the inequality is reversed [^1] In some literature, (2) for $r\geq 1$ is called Lieb–Thirring inequality [4], and the reversed inequality for $r\in[0,1]$ is sometimes called the Araki–Cordes inequality [5, 6].
+In this paper, we will use the term Araki–Lieb–Thirring inequality, referring to the inequality for all $r\geq 0$.
+for $r\in[0,1]$) play important roles in statistical mechanics, quantum information theory, and quantum computing.
 For example, (1) implies that
@@
 where $D_{\alpha}^{*}(\rho\|\sigma)$ is the sandwiched Rényi divergence [8, 9].
-Essentially, Golden–Thompson inequality and Araki–Lieb–Thirrling inequality manifest the noncommutative nature of quantum mechanics and that certain inequalities instead of the equality will be induced in analytical derivations.
-Araki–Lieb–Thirrling trace inequality (2) can be strengthened to the log-majorization relation [5]:
+Essentially, the Golden–Thompson inequality and the Araki–Lieb–Thirring inequality manifest the noncommutative nature of quantum mechanics and show that certain inequalities, instead of equalities, arise in analytical derivations.
+Araki–Lieb–Thirring trace inequality (2) can be strengthened to the log-majorization relation [5]:
 $\left(A^{1/2}BA^{1/2}\right)^{r}\prec_{\log}A^{r/2}B^{r}A^{r/2},\quad r\geq 1$ (5)
@@
 $\prod_{i=1}^{k}\lambda_{i}(A)\leq\prod_{i=1}^{k}\lambda_{i}(B),\quad k=1,2\ldots,n,$ (6)
-and $\det{A}=\det{B}$, where $\lambda_{i}(A)$ is is $i$-th eigenvalue of $A$ arranged in non-increasing order. Likewise, (1) can be strengthened by (5) and the Lie–Trotter formula [10]:
+and $\det{A}=\det{B}$, where $\lambda_{i}(A)$ is the $i$-th eigenvalue of $A$ arranged in non-increasing order. Likewise, (1) can be strengthened by (5) and the Lie–Trotter formula [10]:
 $\operatorname{\mathrm{e}}^{H+K}\prec_{\log}\left(\operatorname{\mathrm{e}}^{pH/2}\operatorname{\mathrm{e}}^{pK}\operatorname{\mathrm{e}}^{pH/2}\right)^{1/p},\quad\forall p>0.$ (7)
-Ando and Hiai later obtained a log-majorization via Löwner–Heinz operator inequality and the antisymmetric tensor power trick (see e.g., [11, Lemma 2.4]), which complements Golden–Thompson type: [10]:
+Ando and Hiai later obtained a log-majorization via Löwner–Heinz operator inequality and the antisymmetric tensor power trick (see e.g., [11, Lemma 2.4]), which complements the Golden–Thompson type [10]:
 $\left[\mathrm{e}^{pK/2}\left(\mathrm{e}^{-pK/2}\mathrm{e}^{pH}\mathrm{e}^{-pK/2}\right)^{\alpha}\mathrm{e}^{pK/2}\right]^{1/p}\prec_{\log}\operatorname{\mathrm{e}}^{\alpha H+(1-\alpha)K},\quad\alpha\in[0,1],\;p>0.$ (8)
-The above then implies that
+The result then implies that
 $D_{\alpha}^{\flat}(\rho\|\sigma)\leq\frac{1}{\alpha-1}\log\operatorname{Tr}\left[\rho\left(\rho^{-1/2}\sigma\rho^{-1/2}\right)^{1-\alpha}\right]=:\widehat{D}_{\alpha}(\rho\|\sigma),\quad\alpha\in(0,1].$ (9)
-Bebiano, Lemos and Providência applied Ando and Hiai’s proof technique with Furuta’s operator inequality [12] to obtain the following log-majorization of Araki’s type.
+In addition, two relatively recent works further generalize these classical results: Bourin and Lee established a family of matrix inequalities from a two-variable functional, which in particular provides an alternative proof of the Golden–Thompson inequality in log-majorization (see [12, Corollary 2.12]), while Hiai developed a generalization of Araki’s log-majorization via a log-convexity theorem [13, Theorem 3.1].
+Bebiano, Lemos, and Providência applied Ando and Hiai’s proof technique, along with Furuta’s operator inequality [14], to obtain the following log-majorization of Araki’s type.
 Theorem BLP (BLP Inequality [15]).
@@
 $\operatorname{Tr}\left[A\left(\log A+\log B\right)\right]\leq\frac{1}{p}\operatorname{Tr}\left[A\log\left(A^{p/2}B^{p}A^{p/2}\right)\right],\quad p>0$ (12)
-(see also the early paper by Hiai and Petz [14], the complementaion by Ando and Hiai [10], and the recent developments by Lieb and Carlen [15]).
+(see also the early paper by Hiai and Petz [16], the complementation by Ando and Hiai [10], and the recent developments by Lieb and Carlen [17]).
 Inequality (12) then implies that the following relation between Umegaki’s quantum relative entropy $D(\rho\|\sigma)$ [18] and Belavkin–Staszewski’s relative entropy $\widehat{D}(\rho\|\sigma)$ [19], i.e.,
 $D(\rho\|\sigma):=\operatorname{Tr}\left[\rho(\log\rho-\log\sigma)\right]\leq\operatorname{Tr}\left[\rho\log\left(\rho^{1/2}\sigma^{-1}\rho^{1/2}\right)\right]=:\widehat{D}(\rho\|\sigma).$ (13)
-With efforts to develop the reversed BLP inequalities [18, 19, 20], the above log-majorizations can be unified as the following generalized BLP (GBLP) inequalities.222 Some literature studied different generalizations of the BLP inequality, (11), and used the term generalized BLP [21, 22, 23, 24].
-In this paper, we call the inequalities and the reversed versions with the possible range of parameters as in Theorem GBLP as the GBLP inequality.
+With efforts to develop the reversed BLP inequalities [20, 21, 22], the above log-majorizations can be unified as the following generalized BLP (GBLP) inequalities.
+[^2] Some literature has studied different generalizations of the BLP inequality, (11), and used the term generalized BLP [23, 24, 25, 26].
+In this paper, we refer to the inequalities and the reversed versions with the possible range of parameters as in Theorem GBLP as the GBLP inequality.
 Theorem GBLP (GBLP Inequality [15, 20, 21, 22]).
@@
 $A^{r+q}B^{q}\prec_{\log}A^{r}\left(A^{p/2}B^{p}A^{p/2}\right)^{q/p},\quad 0<q\leq p,\;r\geq 0.$ (14)
-If either $0\leq r\leq p\leq q$ and $p>0$, or $0\leq q<p$ and $-r\geq q$, then (14) holds with reversed log-majorization.
-Theorem GBLP and together with a result by Matharu and Aujla (obtained similarly from the Furuta inequality) [25] show the following relations:
+If either $0\leq r\leq p\leq q$ and $p>0$, or $0\leq q<p$ and $r\geq q$, then (14) holds with reversed log-majorization.
+Theorem GBLP, together with a result of Matharu and Aujla [27] (which follows from Ando and Hiai’s work [10, Corollary 2.3], or directly from the Furuta inequality [14]), yields the following relations:
 $\begin{split}A\left(A^{-1/2}BA^{-1/2}\right)^{s}&\preceq_{\log}A^{1-s}B^{s},\quad s\in[0,1]\cup[2,\infty);\\ A\left(A^{-1/2}BA^{-1/2}\right)^{s}&\succeq_{\log}A^{1-s}B^{s},\quad s\in[1,2].\end{split}$ (15)
-The relations given in (15) is intriguing in the following sense.
-Considering $s\in[0,1]$, then Araki–Lieb–Thirrling inequality shows that $\left(A^{-1/2}BA^{-1/2}\right)^{s}\succeq_{\log}A^{-s}B^{s}$.
+The relations given in (15) are intriguing in the following sense.
+Considering $s\in[0,1]$, then Araki–Lieb–Thirring inequality shows that $\left(A^{-1/2}BA^{-1/2}\right)^{s}\succeq_{\log}A^{-s}B^{s}$.
 However, when multiplying both sides by $A$, distributing the power $s$ yields a reversed log-majorization $A\left(A^{-1/2}BA^{-1/2}\right)^{s}\preceq_{\log}A^{1-s}B^{s}$.
-Hence, the goal of this paper is to identify which function $f$ such that the following trace inequality holds,
+Hence, the goal of this paper is to identify for which functions $f$ the following trace inequality holds,
 $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],$ (16)
@@
 Our main result is to show that any nonnegative and nondecreasing function $f$ satisfies (16) (see Theorem 1), and, for any $s\in[0,1]$ and any function $g$ such that $x\mapsto x^{s}g(x)$ is nonnegative and nonincreasing, (16) is reversed (see Proposition 2).
-We provide our main results and the proofs in Section 2. In Section 3, we discuss the scneario for $s>1$.
+We provide our main results and the proofs in Section 2. In Section 3, we discuss the scenario for $s>1$.
 2. An Araki-type trace inequality
-We first show that (16) holds if $f(A)$ is a kind of orthogonal projection matrices.
+We first show that (16) holds if $f(A)$ is a specific type of orthogonal projection matrix.
 Proposition 0.
 Let $A=\sum_{i=1}^{n}\lambda_{i}P_{i}$ be the spectral decomposition of positive semi-definite $A$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, and let $\Pi_{k}:=\sum_{i=1}^{k}P_{i}$.
-Then, for all $B\geq 0$,
-[…]
+Then, for all $B\geq 0$, $\operatorname{Tr}\left[\Pi_{k}A^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],\,1\leq k\leq n.$
 Remark 2.1.
-The proposition 1 holds for $\Pi_{k}$ being a projection onto the eigenspace associated with the $k$-th largest eigenvalues.
+Proposition 2 holds for $\Pi_{k}$ being a projection onto the eigenspace associated with the top $k$ eigenvalues.
 The trace inequality may not be true once replacing $\Pi_{k}$ by an arbitrary eigenprojection $P_{i}$.
-To see a counter-example, let $s=1/2$ and
+To see a counterexample, let $s=1/2$ and
 $A=\begin{bmatrix}2&0\\ 0&1\end{bmatrix},\,B=\begin{bmatrix}1&1\\ 1&1\end{bmatrix},\,P_{2}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$. (17)
@@
 First, note that $A$ and each $\tilde{A}_{k}$ commute.
-Let $R=(A\tilde{A}_{k}^{-1})^{1/2}$ be a Hermitian contraction, i.e., $R^{2}\leq I$.
-Since $x\mapsto x^{s}$ is operator concave for $s\in[0,1]$, the operator Jensen’s inequality [26] implies that
-$\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}=R\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}R\leq\left(R\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}R\right)^{s}=\left(A^{1/2}BA^{1/2}\right)^{s}.$
-Hence, we obtain
+Let $R=(A\tilde{A}_{k}^{-1})^{1/2}$, which is a Hermitian contraction, i.e., $R^{2}\leq I$.
+For $s\in(0,1]$, since the function $x\mapsto x^{s}$ is operator concave on $0,\infty)$ and vanishes at $x=0$, the operator Jensen’s inequality [28] implies that
+$\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}=R\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}R\leq\left(R\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}R\right)^{s}=\left(A^{1/2}BA^{1/2}\right)^{s},$
+and the case $s=0$ is trivial. Therefore, we obtain
 $\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right]$ $=\operatorname{Tr}\left[\Pi_{k}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\left(A\tilde{A}_{k}^{-1}\right)^{1/2}\right]$ (18)
 $\leq\operatorname{Tr}\left[\Pi_{k}\left(A^{1/2}BA^{1/2}\right)^{s}\right],$ (19)
-with the fact that $\Pi_{k}(A\tilde{A}_{k}^{-1})^{1/2}=(A\tilde{A}_{k}^{-1})^{1/2}\Pi_{k}=\Pi_{k}$.
-The case of $s\in[1,2]$ follows similarly by the operator convexity of $x\mapsto x^{s}$, $s\in[1,2]$. ∎
+using the fact that $\Pi_{k}(A\tilde{A}_{k}^{-1})^{1/2}=(A\tilde{A}_{k}^{-1})^{1/2}\Pi_{k}=\Pi_{k}$.
+The case $s\in[1,2]$ follows similarly, using instead the operator convexity of $x\mapsto x^{s}$, $s\in[1,2]$. ∎
 Note that the BLP inequality implies that, for $A,B\geq 0$,
@@
 $\operatorname{Tr}\left[A^{t}A^{s}\left(A^{-1/2}BA^{-1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}B^{s}\right],\quad\forall\,s\in[0,1],\,t\geq 0.$
-Next, substitute $A$ by $A^{-1}$, then
+Next, substitute AA by $A^{-1}$, then
 $\operatorname{Tr}\left[A^{-t-s}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{-t}B^{s}\right],\quad\forall\,s\in[0,1],\;t\geq 0.$
 Or,
-$\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right],\quad\forall\,s\in[0,1],\;t\leq-s.$
+$\operatorname{Tr}\left[A^{t}\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[A^{t}A^{s}B^{s}\right],\quad\forall\,s\in[0,1],\;t\leq-s.\qed$
 Proof of Proposition 2.
-Combining the two different conditions regarding $t$ in (20) and Lemma 3, we conclude that for $f(x)=x^{t}$, $t\geq 0$ or $f(x)=-x^{t}$, $t\leq-s$, we have
+Combining the two different conditions regarding $t$ from (20) and Lemma 2, we conclude that for $f(x)=x^{t}$, $t\geq 0$ or $f(x)=-x^{t}$, $t\leq-s$, we have
 $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$ (21)
@@
 $\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{s}B^{s}\right]\leq\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$ (22)
-To show this, we first introduce $\tilde{A}_{k,\varepsilon}=\Pi_{k}A+(\lambda_{k}-\varepsilon)\Pi_{k}^{\mathrm{c}}$ as an approximation of $\tilde{A}_{k}$, for $\varepsilon\in(0,\lambda_{k})$, where $\lambda_{k}$ is $k$-th largest eigenvalue of $A$.
+To show this, we first introduce $\tilde{A}_{k,\varepsilon}=\Pi_{k}A+(\lambda_{k}-\varepsilon)\Pi_{k}^{\mathrm{c}}$ as an approximation of $\tilde{A}_{k}$, for $\varepsilon\in(0,\lambda_{k})$, where $\lambda_{k}$ is the $k$-th largest eigenvalue of $A$.
 Define
@@
 $\operatorname{Tr}\left[g_{t}(\tilde{A}_{k,\varepsilon})\tilde{A}_{k,\varepsilon}^{s}B^{s}\right]\leq\operatorname{Tr}\left[g_{t}(\tilde{A}_{k,\varepsilon})\left(\tilde{A}_{k,\varepsilon}^{1/2}B\tilde{A}_{k,\varepsilon}^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1],\,t\leq-s,$
-with noticing that the constant function $1$ is just $f(x)=x^{t}$ at $t=0$.
+noting that the constant function $1$ is just $f(x)=x^{t}$ at $t=0$.
 Let $t\to$-$\infty$, then $g_{t}(\tilde{A}_{k,\varepsilon})\to\Pi_{k}$, and hence
@@
 Theorem 1.
-Let $A,B\geq 0$. For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
-$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
+Let $A,B\geq 0$.
+For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\leq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\in[0,1].$
 Proof.
@@
 Let $A,B\geq 0$.
-For any $s\in[0,1]$ and any function $g$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, such that $x\mapsto x^{s}g(x)$ is nonnegative and nonincreasing, then
-$\operatorname{Tr}\left[g(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].$
+For any $s\in[0,1]$ and any function $g$ on an interval $\mathcal{J}\subseteq[0,\infty)$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, such that $x\mapsto x^{s}g(x)$ is nonnegative and nonincreasing, then $\operatorname{Tr}\left[g(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right]\leq\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].$
 Proof.
-Again, we assume $A>0$, and extend to $A\geq 0$ via a limit.
+Again, we assume $A>0$, and extend the result to $A\geq 0$ via a limit.
 Set $f(x)=x^{-s}g(x^{-1})$, which is nonnegative and nondecreasing.
@@
 $\leq\operatorname{Tr}\left[f\left(A^{-1}\right)\left(A^{-1/2}\cdot A^{{1}/{2}}BA^{{1}/{2}}\cdot A^{-1/2}\right)^{s}\right]$ (24)
-$=\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].$ (25)
+$=\operatorname{Tr}\left[g(A)A^{s}B^{s}\right].\qed$ (25)
 3. A Conjecture for $s>1$
@@
 For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
-$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s>1.$
-Note that $x\mapsto x^{s}$ is operator convex for $s\in[1,2]$, we may adapt the previous technique to prove Conjecture 1 for $s\in[1,2]$, provided that the condition $0\leq r\leq p\leq q$ in the reverse direction of Theorem GBLP can be relaxed to $0\leq p\leq q$.
-Thereby, we conjecture the following reverse log-majorization, which complements (14).
+$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad\forall\,s\geq 1.$
+Note that $x\mapsto x^{s}$ is operator convex for $s\in[1,2]$.
+We may therefore adapt the previous technique to prove Conjecture 1 for $s\in[1,2]$, provided that the condition $0\leq r\leq p\leq q$, $p>0$ in the reverse direction of Theorem GBLP can be relaxed to $0<p\leq q$, $r\geq 0$.
+Therefore, we conjecture the following reverse log-majorization, which complements (14).
 Conjecture 2.
 For $A,B\geq 0$,
-$\operatorname{Tr}\left[A^{r+q}B^{q}\right]\geq\operatorname{Tr}\left[A^{r}\left(A^{p/2}B^{p}A^{p/2}\right)^{q/p}\right],\quad 0\leq p\leq q,\;r\geq 0.$ (26)
+$A^{r+q}B^{q}\succ_{\log}A^{r}\left(A^{p/2}B^{p}A^{p/2}\right)^{q/p},\quad 0<p\leq q,\;r\geq 0.$ (26)
 (To show Conjecture 1 for $s\in[1,2]$, the trace inequality version of Conjecture 2 would be sufficient.)
@@
 Proposition 1.
-For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have
-$\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad s=2.$
+For any nonnegative and nondecreasing function $f$ on an interval $\mathcal{J}$, where $\operatorname{spec}(A)\subseteq\mathcal{J}$, we have $\operatorname{Tr}\left[f(A)A^{s}B^{s}\right]\geq\operatorname{Tr}\left[f(A)\left(A^{1/2}BA^{1/2}\right)^{s}\right],\quad s=2.$
 Proof.
@@
 Lemma 1.
-Let $A=\sum_{i=1}^{n}\lambda_{i}P_{i}$ be the spectral decomposition of positive semi-definite $A$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d}\geq 0$, $\Pi_{k}:=\sum_{i=1}^{k}P_{i}$, $\Pi_{k}^{\textnormal{c}}:=I-\Pi_{k}$, and $\tilde{A}_{k}:=\Pi_{k}A+\lambda_{k}\Pi_{k}^{\textnormal{c}}$.
+Let $A=\sum_{i=1}^{n}\lambda_{i}P_{i}$ be the spectral decomposition of positive semi-definite $A$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, $\Pi_{k}:=\sum_{i=1}^{k}P_{i}$, $\Pi_{k}^{\textnormal{c}}:=I-\Pi_{k}$, and $\tilde{A}_{k}:=\Pi_{k}A+\lambda_{k}\Pi_{k}^{\textnormal{c}}$.
 Then,
@@
 Proof.
-Using the fact that $\lambda_{k}\Pi_{k}\leq\Pi_{k}\tilde{A}_{k}\Pi_{k}$ by construction and the Araki–Lieb–Thirrling trace inequality (2), we calculate
+Using the fact that $\lambda_{k}\Pi_{k}\leq\Pi_{k}\tilde{A}_{k}\Pi_{k}$ by construction and the Araki–Lieb–Thirring trace inequality (2), we calculate
 $\operatorname{Tr}\left[\Pi_{k}\left(\tilde{A}_{k}^{1/2}B\tilde{A}_{k}^{1/2}\right)^{2}\right]$ $=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}\Pi_{k}B{\tilde{A}_{k}}B\right]$ (28)
@@
 $=\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k}\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (34)
-(2)${\leq}\operatorname{Tr}\left[(\Pi_{k}\tilde{A}_{k}\Pi_{k})^{2}(\Pi_{k}B\Pi_{k})^{2}\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (35)
+[…]
 $=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B\Pi_{k}B\right]+\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B{\Pi_{k}^{\text{c}}}B\right]$ (36)
 $=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}\Pi_{k}B(\Pi_{k}+{\Pi_{k}^{\text{c}}})B\right]$ (37)
-$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}B^{2}\right].$ (38)
+$=\operatorname{Tr}\left[\Pi_{k}\tilde{A}_{k}^{2}B^{2}\right].\qed$ (38)
 4. Conclusions
-In this paper, we derived some Araki-type trace inequalities that involve monotone functions other than powers.
-Although usual Araki-type inequalities and the BLP inequalities are expressed in a stronger form of log-majorization, however, the antisymmetric tensor power trick may not be applicable for functions beyond powers.
-On the other hand, trace inequalities find more and immediate applications in other fields such as quantum information science and mathematical physics, which motivates us to further Araki-type trace inequalities to a broader form.
-Numerical experiments suggest that Conjecture 1 for $s>1$ and a slightly general reverse BLP inequality in Conjecture 2 could be true, which is left for future work.
+In this paper, we have derived Araki-type trace inequalities that involve general monotone functions, extending beyond simple powers.
+Although usual Araki-type inequalities and the BLP inequalities are expressed in a stronger form of log-majorization, the antisymmetric tensor power trick may not be applicable to this broader class of functions.
+On the other hand, trace inequalities have broad and direct applications in fields such as quantum information science and mathematical physics.
+This motivates our effort to extend Araki-type trace inequalities to a more general setting.
+Numerical experiments suggest that Conjecture 1 for $s>1$ and a slightly generalized reverse BLP inequality in Conjecture 2 could be true, which remains for future work.
+Acknowledgments
+We are supported by the Emerging Young Scholars Program of the National Science and Technology Council, Taiwan (R.O.C.) under Grants No.
+NSTC 114-2628-E-002 -006, NSTC 114-2119-M-001-002, and NSTC 114-2124-M-002-003, by the Yushan Young Scholar Program of the Ministry of Education, Taiwan (R.O.C.) under Grants No.
+NTU-114V2016-1 and by the research project ‘Forefront Quantum Computing, Learning, and Engineering in Noisy Intermediate-Scale Quantum Era’ of National Taiwan University under Grant NTU-114L895005.
+H.-C.
+Cheng acknowledges the support from the ‘Center for Advanced Computing and Imaging in Biomedicine (NTU-114L900702)’ through The Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE) in Taiwan.